OpenStudy (anonymous):

determine the derivative y' at the point (1,0) y=ln(x^2+y^2)

myininaya (myininaya):

[ln(inside)]'=(inside)'/inside

OpenStudy (amistre64):

y' = (2x + 2yy')/(x^2+y^2) right?

myininaya (myininaya):

yes!

myininaya (myininaya):

solve for y'

OpenStudy (amistre64):

y' = 2x/(x^2 +y^2 - 2x) .. if i did it right...which is a big if :)

OpenStudy (amistre64):

that last 2x should be a 2y...

OpenStudy (amistre64):

y' = 2x/(x^2 + y^2 - 2y) (x=1,y=0) y' = 2/(1) = 2 ??